OpenStudy (anonymous):

find the domain and range of the function: f(g(x)) With the equations: f(x)=(x^2)-6 g(x)=2x-3

OpenStudy (abmon98):

(2x-3)^2-6 domain is 3/2 and the range is -6

OpenStudy (anonymous):

f(g(x))= (2x-3)^2 - 6 4x^2 +9-12x-6 4x^2 -12x +3 So, domain is all the values of x such that \[x \in R \] RANGE: first find the point with slope=0 so differentiating wrt x 8x-12=0 x=3/2 putting the value of x f(g(x))=4(3/2)^2 - 12(3/2) + 3 =-6 so, the range is [-6 , infinity]